A real estate investment has the following expected cash flows:

Year Cash Flows

1 $10,000

2 25,000

3 50,000

4 35,000

The initial cost is $85,000. The discount rate is 8.25 percent. What is the investment's present value?

a). $10,479

b). $13,479

c.) $16,479

d). $14,479

e). $15,479
Answer #1
Year Cash Flow
0 -85000
1 10000
2 25000
3 50000
4 35000

Initial cost = C0 = -85000

Cash flow in year 1 = C1 = 10000

Cash flow in year 2 = C2 = 25000

Cash flow in year 3 = C3 = 50000

Cash flow in year 4 = C4 = 35000

Discount rate = r = 8.25%

Present value is calculated using the formula:

Present Value of the cash flows = PV = C0 + C1/(1+r)1 + C1/(1+r)2 + C1/(1+r)3 + C1/(1+r)4 = -85000 + 10000/(1+8.25%)1 + 25000/(1+8.25%)2 + 50000/(1+8.25%)3 + 35000/(1+8.25%)4 = -85000 + 9237.8752886836 + 21334.5849623178 + 39417.2470435433 + 25489.212868804 = 10478.9201633486 ~ $10479 (Rounded to nearest dollar)

Answer -> $10,479 (Option a)